OpenStudy (anonymous):

What is 1337% of 3.14?

OpenStudy (mr.math):

\[1337\%=13.37\]

OpenStudy (mr.math):

Now, you just need to multiply this number by 3.14 and that will give you 1337% of 3.14.

OpenStudy (anonymous):

Its 42 :O

OpenStudy (mr.math):

Can you multiply 13.37 times 13.14?

OpenStudy (anonymous):

Yeah.. Why?

OpenStudy (anonymous):

Roughly 42..

OpenStudy (mr.math):

Because that's what you're looking for.

OpenStudy (anonymous):

Thanks :D

OpenStudy (mr.math):

That's your answer, read what I wrote above again

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

Whatever, ill go back to my algebra

OpenStudy (anonymous):

btw its 175

OpenStudy (mr.math):

What is 175?

OpenStudy (anonymous):

13.37*13.14

OpenStudy (mr.math):

It's not!

OpenStudy (mr.math):

I meant 3.14 :D

OpenStudy (anonymous):

41.98 :D

OpenStudy (mr.math):

That was a typo.

OpenStudy (anonymous):

I thought it would form a super awesome number and it did!

OpenStudy (mr.math):

lol

OpenStudy (anonymous):

Havent you seen that movie? :D

OpenStudy (mr.math):

Which movie?

OpenStudy (anonymous):

Cant really remember what its called but according to it 42 is the meaning of life xd

hero (hero):

I somehow made a multiplication mistake using my calculator. Here's the correct solution: is/of = %/100 x/3.14 = 1337/100 100x = 4198.18 x = 41.98 which is approximately 42

OpenStudy (jamesj):

"Hitchhikers Guide to the Galaxy". In this case, it really is true that the book is better!